Solve the following system of 3 equations 2x+4y-z=1 and 3x+3y+z=1 and -5x-8y+6z=5
Answer by Edwin McCravy(20056)

By what method? There are several different methods.  Tell me
which method in case you have to show your work.  Tell me in
the note form below and I'll get back to you by email.  No
charge.

The answer is x = -45/31, y = 41/31, z = 43/31

Edwin